OpenStudy (anonymous):

f(x+h)-f(x) / h for f(x) = 3x^2 - x - 6

OpenStudy (anonymous):

\[\frac{ f(x+h)-f(x) }{ h }\] \[f(x) = 3x ^{2}-x-6\]

OpenStudy (anonymous):

If it is just that function, then you can plug in (x+h) in the function F(x) and then simply subtract the equation f(x). This looks like a derivative problem, so if there is a \[\lim_{h \rightarrow 0}\] before that whole function, then the answer would be the derivative of the function f(x), which in this case is: \[6x-1\] If there is no limit, then you wind up with: \[\frac{[(x+h)^2-(x+h)-6]- (3x^2-x-6)}{ h }\] Which you can simplify by multiplying out the terms and adding like terms.

OpenStudy (anonymous):

Do you mean: \[\frac{ (3)(x+h)^{2}-(x+h)-6) - (3x ^{2}-x-6) }{ h }\]

OpenStudy (anonymous):

Or the 3 cannot be included?

OpenStudy (anonymous):

Yes you're right, I missed the three. Include it.

OpenStudy (anonymous):

Can you help me solve. Lol. Im here having problem because Im getting a square.

OpenStudy (anonymous):

So far, this is where I'm at:

OpenStudy (anonymous):

\[\frac{ 3(x+h)^{2}-(x+h)-6-(3x ^{2}-x-6 }{ h }\] So these will cancel out: \[3x^{2}, -3x ^{2}, -x, -x, -6, -6\] Leaving: \[\frac{ 3h ^{2}+h }{ h }\] How did I get \[3h ^{2}? --- 3(x+h)^{2} = 3x+3h ^{2}\]

OpenStudy (ranga):

3(x+h)^2 - (x+h) - 6 = 3(x^2+2xh+h^2) - x - h - 6 = 3x^2 + 6xh + 3h^2 - x -h - 6 subtract 3x^2 - x - 6: 3x^2 + 6xh + 3h^2 - x -h - 6 -3x^2 + x + 6 -------------------------------- 6xh + 3h^2 -h = h(3h + 6x - 1)

OpenStudy (ranga):

Divide by h, and it will cancel out the h in the numerator leaving you with: (3h + 6x - 1)